### v5

# Steering Language Models for Theorem Proving

Abstract
Recent advances in automated theorem proving use Large Language Models (LLMs) to translate informal mathematical statements into formal proofs. However, informal cues are often ambiguous or lack strict logical structure, making it hard for models to interpret them precisely. While existing methods achieve strong performance, little is known about how LLMs internally represent informal cues, or how these influence proof generation. To address this, we explore activation steering, an inference-time intervention that identifies linear directions in residual activations associated with informal reasoning traces and adjusts them to improve proof construction without fine-tuning. This mechanism also yields interpretable information about how reasoning is internally encoded in the activation space of LLMs. We test our method for generating formal proofs from already-formalized theorems. Our contributions are twofold: (1) a novel activation-based intervention for guiding proof synthesis in LLMs; and (2) demonstration that this intervention improves performance under two decoding strategies (sampling and best-first search) without any further training.
1 Introduction
Interactive proof assistants such as Lean (de Moura et al., 2015), Isabelle (Wenzel et al., 2008), and Rocq (Barras et al., 1999) provide the infrastructure for formal verification of mathematical proofs and software. They require proofs to be expressed in precise formal languages (Avigad, 2023; Ringer et al., 2019). Neural theorem proving, combining LLMs with proof assistants, has shown promise in automating reasoning (First et al., 2023; Polu & Sutskever, 2020; Polu et al., 2022; Yang et al., 2023; Welleck, 2023). Prior work (Lin et al., 2024; Welleck et al., 2021; 2022) suggests that training model on natural reasoning for proof steps can improve performance, but it remains unclear how such informal language is internally represented or whether it can be reliably leveraged to improve proof generation.
Building on the observation that proofs often interleave natural-language reasoning with formal steps, we show that informal natural language (NL) context induces distinct activation patterns in LLMs. We extract these patterns as steering vectors as in (Panickssery et al., 2024b; Turner et al., 2024; Lucchetti & Guha, 2024), and use them to intervene at inference time. These interventions improve proof generation quality, while preserving model behavior in unrelated aspects. We test our approach across three theorem-proving oriented LLMs - Lemma (Azerbayev et al., 2024), InternLM-2 (Ying et al., 2024), and InternLM2.5-StepProver (Wu et al., 2024). On established benchmarks MiniF2F and PutnamBench, we find that steering via informal language context consistently improves proof quality under multiple decoding strategies.
Our key contributions are threefold: (1) We provide mechanistic insights into how informal mathematical or natural language context impacts internal reasoning in LLMs for theorem proving. (2) We propose a method to extract activation vectors that encode such informal context, and use these vectors to guide proof construction in a structured, grounded way. (3) We demonstrate that activation steering yields consistent improvements in proof success rates on MiniF2F and PutnamBench benchmarks under both search- and sampling-based decoding.
By focusing on steering in activation space, our approach sheds light on the link between informal mathematical language and formal reasoning steps; without needing to fine-tune model weights. We base this on the hypothesis that many reasoning features are encoded as (approximately) linear directions in activation space (Mikolov et al., 2013; Elhage et al., 2021; Nanda et al., 2023; Park et al., 2024). Although not all features follow perfect linearity (Engels et al., 2025), this assumption has previously enabled methods like concept erasure and steering to work well (Beaglehole et al., 2025; Zhao et al., 2025a; Shah et al., 2025).
The rest of this paper is organized as follows. In Section 2, we review related work on neural theorem proving, activation steering, and the representation of informal mathematical reasoning in LLMs. Section 3 presents our activation steering method: how steering vectors are constructed from informal mathematical contexts, how they are applied at inference time, and how we select layers and intervention strengths. Section 5 describes our experimental setup, including models, benchmarks (MiniF2F, PutnamBench), and decoding strategies (sampling vs best-first search). Section 6 reports results: we analyze performance gains, the effect of steering vectors on proof success rates, and ablations experiments. Finally, in Section 7 we discuss insights into the internal reasoning of the model, limitations of our approach, and potential directions for future work.
2 Related Work
This section situates our work in three intersecting strands: automatic theorem proving and auto‐formalization, representation learning in large language models (LLMs), and mechanistic interpretability via activation interventions.
Automatic Theorem Proving and Auto-formalization Recent progress in automatic theorem proving frames proof search as sequence generation. The GPT-f framework (Polu & Sutskever, 2020) trains a language model to map proof states to tactics, combining it with best‐first search to assemble proofs. Extensions explore data augmentation via proof transformations or synthesis (Han et al., 2022; Rotella et al., 2025; Wang & Deng, 2020), improved search strategies (Wang et al., 2023), curriculum training (Polu et al., 2022), and retrieval from proof libraries (Yang et al., 2023). Systems like LLMStep unify these ideas into usable frameworks (Welleck & Saha, 2023).
Auto‐formalization complements this by translating informal mathematics (e.g. text, sketches) into formal proofs. Surveys highlight translation techniques (Wu et al., 2022), while Draft-Sketch-Prove shows LLMs prompted with informal sketches outperform formal-only baselines (Jiang et al., 2023). LeanStar (Lin et al., 2024) interleaves informal reasoning with tactic prediction, typically via supervised fine-tuning on synthetic data. A central open question remains: How do informal reasoning patterns inform formal proving within a model’s internal representations? Our approach explores this via steering vectors in activation space, injecting implicit natural-language “thoughts” at inference to guide proof steps without heavy fine-tuning, probing the latent link between informal intuition and formal reasoning.
Language Model Representation of Concepts Our method is motivated by prior work demonstrating that task or feature concepts can be represented as linear directions in the activation space of LLMs when given appropriate contextual examples. Hendel et al. (2023) and Todd et al. (2024) show that a context (e.g. a prompt) can induce a task embedding in some activation subspace. A broader literature examines linear decompositions of features such as truthfulness (Azaria & Mitchell, 2023; Li et al., 2023; Marks & Tegmark, 2024), sentiment (Tigges et al., 2024), harmlessness (Zou et al., 2025; Zheng et al., 2024), sycophancy or alignment steering (Perez et al., 2023; Panickssery et al., 2024a; Sharma et al., 2024), factual knowledge (Gurnee & Tegmark, 2024), and refusal behavior (Arditi et al., 2024). Relatedly, unsupervised methods like sparse autoencoders have been used to extract concept directions in hidden space (Bricken et al., 2023; Huben et al., 2024; Templeton et al., 2024). These works generally share the hypothesis that LLMs encode high-level features or concepts as (approximately) linear directions in activation space (Elhage et al., 2021; Mikolov et al., 2013; Nanda et al., 2023; Park et al., 2024). While recent work cautions that not all features may admit clean linear representations (Engels et al., 2025), the linearity assumption has proven effective in practice for concept erasure, model steering, and interpretability (Beaglehole et al., 2025; Zhao et al., 2025a; Shah et al., 2025). Within this framework, we explore whether generating a natural-language informal explanation can itself be represented as a linear direction, and how injecting this direction at inference improves formal theorem proving performance.
Mechanistic Interpretability and Activation Patching A rich body of work in mechanistic interpretability seeks to locate, analyze, and manipulate internal representations in transformer-based models. Early studies localized factual knowledge or associative memory to particular neurons or circuits (Meng et al., 2022), and probed hidden layers for high-level features (Li et al., 2024; Dong et al., 2023). The idea of implicit evaluation i.e. measuring latent capability of a model rather than just output behavior, has been developed to complement benchmark-based evaluation (Dong et al., 2023). One particularly relevant tool is activation patching (or residual stream intervention) (Vig et al., 2020; Variengien & Winsor, 2023), wherein one alters specific activations in a layer (often in the residual stream) at inference time to influence model outputs. Recent works have used this to edit factual associations or steer behavior (Zhao et al., 2025b). The residual stream is a high-dimensional accumulator of intermediate features (propagated via skip connections) that each transformer layer refines or adds to; intervening there offers a principled way to steer model behavior (Elhage et al., 2021). In the domain of theorem proving, activation interventions provide a promising lens into how a model processes and integrates informal guidance with formal reasoning. In our approach, we generate steering vectors (patches) that, when applied to the residual stream at certain layers, help the model emit an internal “natural-language thought” prior to or alongside tactic prediction. We find that effective steering tends to realign activations such that the output of the model conforms to a structured reasoning format, thereby improving downstream proof search and reducing failure modes.
Building on these insights, we adapt activation interventions to theorem proving by treating informal reasoning as a direction in activation space. We then show how extracting and injecting this direction can guide proof generation. The next section details the architecture, design choices, and evaluation of this steering approach.
3 Steering for improving theorem proving
We aim to steer theorem-proving models toward using informal reasoning via activation interventions. This section describes (i) how we compute steering vectors, (ii) how we choose layers and apply steering, and (iii) efficiency and practical considerations.
3.1 Intuition and Overview
Modern transformer models often encode high-level semantic behavior (e.g. reasoning, style) as approximately linear directions in activation space (residual streams) (Zou et al., 2025; Elhage et al., 2021). We exploit this property: given pairs of prompts that differ only by the presence of natural-language reasoning (but otherwise describe the same proof step), we can estimate a vector in activation space that points in the “informal reasoning” direction. At inference time, adding this vector (appropriately scaled) nudges the model toward producing reasoning-augmented proofs.
In the rest of this section, we detail how we compute these steering vectors, how we pick which layers to intervene on, and how we incorporate them efficiently at inference.
3.2 Constructing Steering Vectors
We adopt a difference-of-means (contrastive) method (Belrose et al., 2023), which has been effective at extracting feature directions across multiple domains (e.g. refusal, truthfulness) (Arditi et al., 2024; Panickssery et al., 2024b; Marks & Tegmark, 2024). Let $\mathcal{D}=\{(p_{i},p_{i}^{+})\}_{i=1}^{N}$ be a dataset of paired prompts, where $p_{i}$ is a standard formal proof prompt and $p_{i}^{+}$ is its version augmented with explicit natural-language reasoning. We feed both prompts into our model $\mathcal{M}$ and extract the residual stream activations at a chosen layer $\ell$. Denote the activation at the final token position by $\mathbf{v}_{\ell}^{p}=\text{resid}(\mathcal{M}(p),\ell)$.
Then the steering vector $\mathbf{u}^{(\ell)}$ is computed as:
$\mathbf{u}^{(\ell)}=\frac{1}{|\mathcal{D}^{+}|}\sum_{p^{+}\in\mathcal{D}^{+}}\mathbf{v}_{\ell}^{p^{+}}-\frac{1}{|\mathcal{D}^{-}|}\sum_{p^{-}\in\mathcal{D}^{-}}\mathbf{v}_{\ell}^{p^{-}}$ (1)
Here $\mathcal{D}^{+}$ and $\mathcal{D}^{-}$ are the two halves of the prompt-pair dataset (augmented and unaugmented), so the subtraction isolates the direction most correlated with natural-language reasoning while largely cancelling out shared biases or irrelevant activation patterns.
Figure 1: Steering Vectors are computed as difference of means for $p^{+}$ and $p$
We emphasize that this method requires only forward passes; no gradient-based finetuning or parameter updates; making it computationally lightweight.
3.3 Layer Selection and Intervention Strategy
Not all layers are equally responsive to steering, so we first perform an activation analysis to find suitable intervention points. For each layer $\ell$, we measure:
$\text{sim}(\ell)=\frac{1}{|\mathcal{D}|}\sum_{(p,p^{+})\in\mathcal{D}}\cos(\mathbf{v}_{\ell}^{p},\mathbf{v}_{\ell}^{p^{+}})$ (2)
Empirically, we observe that in early layers, $\text{sim}(\ell)$ is close to 1 (i.e. little divergence), but in deeper layers it drops and exhibits local minima (“valleys”) as seen in Figure 2.
Figure 2: Cosine similarity across model’s residual stream activations for each layer.
We find that in early layers, activations remain highly similar between $p$ and $p^{+}$, but as we go deeper, the cosine similarity drops and exhibits local minima. Intuitively, these are the layers where the representations diverge most when informal reasoning is introduced (see Figure 2). Thus, these “valley” layers are promising candidates for steering. These valleys likely correspond to points where the internal representation of the model is most sensitive to reasoning-specific perturbations.
We select intervention layers where (a) cosine similarity dips significantly, and (b) representation is still semantically stable. At inference, we inject the steering vector into those layers:
$\mathbf{v}_{\ell}^{\prime}=\mathbf{v}_{\ell}+\alpha\cdot\mathbf{u}^{(\ell)}$ (3)
We treat $\alpha$ as a hyperparameter and specify that it will be tuned on held-out validation data. The method instantiates a sweep to find the optimal balance between influencing reasoning and maintaining validity. We place constraints (e.g. an upper bound on $\alpha$) to avoid destabilizing proofs. We present the complete pseudocode of our approach in Algorithm 1.
Algorithm 1 Activation Steering for Theorem Proving
0: Contrastive prompt pairs $D=\{(p_{i},p_{i}^{+})\}_{i=1}^{N}$ where $p_{i}$ are standard prompts and $p_{i}^{+}$ include natural language reasoning
0: Language model $M$ with $L$ layers
0: Test theorem $T$ to prove
0: Enhanced proof generation
1:Offline: Create Steering Vectors
2:for each layer $\ell\in\{1,\ldots,L\}$ do
3:  $\bar{v}^{+}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\text{resid}(M(p_{i}^{+}),\ell)$ {Mean activation for augmented prompts}
4:  $\bar{v}^{-}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\text{resid}(M(p_{i}),\ell)$ {Mean activation for standard prompts}
5:  $u^{(\ell)}\leftarrow\bar{v}^{+}-\bar{v}^{-}$ {Steering vector at layer $\ell$}
6:end for
[…]
8:Online: Apply Steering During Inference
9: Initialize proof context with theorem $T$
10:while proof incomplete and within computational budget do
11:  Perform forward pass through model
12:  for each intervention layer $\ell\in\mathcal{L}$ do
13:   $v^{\prime}_{\ell}\leftarrow v_{\ell}+\alpha\cdot u^{(\ell)}$ {Inject steering vector}
14:  end for
15:  Generate next proof step using modified activations
16:  Validate and apply proof step if correct
17:end while
18:return Generated proof or failure
4 Experimental Setup
4.1 Models
We conduct experiments on three open-source 7B-parameter language models optimized for mathematical reasoning:
  • Llemma-7B (Azerbayev et al., 2024): Built on Code Llama architecture, pre-trained on mathematical texts and formal mathematics corpora.
  • InternLM2-7B (Ying et al., 2024): LLaMA-based architecture with extended training on mathematical problem-solving.
  • InternLM2.5-StepProver (Wu et al., 2024): Enhanced variant with expert iteration on large-scale Lean problems.
All models share transformer architectures with consistent residual stream dimensionality, facilitating uniform steering vector extraction.
4.2 Data and Vector Construction
We construct steering vectors from the Lean-STaR dataset(Lin et al., 2024), randomly sampling 10,000 theorem-proof pairs. Each datapoint $(p,p^{+})$ consists of a formal proof step $p$ and its augmented version $p^{+}$ containing explicit natural language reasoning.
For robustness, we generate model responses $r_{p}$ and $r_{p^{+}}$ for each prompt pair and retain only instances where both responses are valid proof steps with $r_{p}\neq r_{p^{+}}$. This filtering yields approximately 7,400 high-quality contrastive pairs.
Vector construction requires only forward passes through the model, consuming $\sim 1$ GPU-hour on NVIDIA A100, a 60× reduction compared to LoRA fine-tuning while achieving superior performance gains.
4.3 Chosen Layers and $\alpha$
Based on activation similarity analysis (Section 3.3), we apply steering vectors to Layers 22, 25 in InternLM2-7B and Layers 24, 25 in Llemma-7B. These correspond to layers where the model’s representation diverges most from baseline while maintaining semantic coherence. We empirically determine $\alpha=0.8$ through validation experiments.
4.4 Evaluation Benchmarks
miniF2F Our primary evaluation utilizes miniF2F (Zheng et al., 2022), a standardized benchmark comprising 244 theorems drawn from mathematical competitions (AMC, AIME, IMO). These problems span algebra, number theory, geometry, and combinatorics with varying proof complexities. We employ two proof search strategies:
  • Best-first search: Expansion budget $N\in\{50,600\}$, tactic sampling width $S=32$, selecting states by cumulative log-probability
  • Parallel sampling: $K$ independent proof attempts to accommodate probability shifts from natural language injection
PutnamBench We additionally evaluate on PutnamBench (Tsoukalas et al., 2024), containing problems from the William Lowell Putnam Mathematical Competition. Following standard evaluation protocol, we test on both the Lean subset (657 problems) and Rocq subset (412 problems) using the benchmark’s default search parameters: $N=600$ expansion budget with $S=32$ tactic width for best-first search, and parallel sampling with $K=2$ attempts. This benchmark features more challenging problems with longer average proof lengths, providing a stringent test of steering effectiveness on complex mathematical reasoning.
Evaluation Metrics.  Following established practice in neural theorem proving, we report pass rates (percentage of theorems successfully proved within computational budget) as our primary metric. For detailed analysis, we additionally examine proof characteristics including average length, tactic distribution, and the frequency of intermediate lemma usage (via the have tactic) to understand how steering affects proof structure and strategy.
5 Results and Analysis
Our experiments were designed to address the following research questions:
  1. 1.
    Does activation steering improve theorem-proving performance of existing models?
  2. 2.
    Which layers contribute most effectively when steered?
  3. 3.
    Does steering enhance proof structure and search efficiency?
  4. 4.
    Are improvements consistent across different search budgets?
  5. 5.
    How effective and efficient are steering vectors compared to LoRA fine-tuning?
  6. 6.
    Do steering vectors generalize across provers?
5.1 Q1: Impact of Activation Steering
Activation steering consistently improves model performance across all tested configurations, enabling the discovery of proofs not derivable by the base models alone.
Model MiniF2F
Llemma 7B 26.6%
InternLM2-7B 28.7%
InternLM2.5-Step Prover 48.2%
LLEMMA-7B + Steering 28.1%
InternLM2-7B + Steering 32.4%
InternLM2.5-StepProver + Steering 66.4%
Table 1: Performance on MiniF2F (sampling decoding, 50×32×1).
Model PutnamBench
InternLM2 7B 4 (0.6%)
InternLM2.5-StepProver 6 (0.9%)
InternLM2 7B + Steering 4 (0.6%)
InternLM2.5-StepProver + Steering 7 (1.1%)
Table 2: Performance on PutnamBench (Lean). Results reported out of 657 attempts.
MiniF2F. As shown in Table 2, steering improves performance across all models on the miniF2F-Test benchmark (Zheng et al., 2022). We adopt the sampling decoding strategy from LeanStar (Lin et al., 2024), which mitigates variance from natural-language generation and provides more effective node selection than standard Best-First Tree Search (see Appendix for Best-First results).
Steering introduces structured natural-language comments that enhance mathematical reasoning and proof generation. Importantly, we find that steering enables a significant number of new proofs beyond the base model’s reach, suggesting that steering vectors guide the model toward otherwise inaccessible solution paths. Interestingly, models occasionally succeed despite incorrect informal reasoning, implying that steering exerts influence beyond surface-level natural language.
While steering increases successful proofs, it also introduces more failure cases. We hypothesize this stems from biases in the data used to construct steering vectors, which encode inductive priors that align well with some theorem classes but misalign with others. Despite this trade-off, the net effect is strongly positive, establishing activation steering as a lightweight, parameter-free alternative to fine-tuning.
PutnamBench. On PutnamBench (Table 2), steering improves success rates for both InternLM2 and InternLM2.5 on Lean, and achieves non-trivial gains on Rocq problems (details in Section A.3). These results highlight steering’s ability to support reasoning in more complex, multi-step proofs.
5.2 Q2: Layer Choice
We evaluate layer sensitivity following the strategy described in Section 3.3. Steering vectors are patched layer-wise to analyze their influence. Figure 4 and Table 4 show that later layers exert stronger influence on theorem-proving performance.
Selected Layers Pass Rate (%)
25–30 (Late) 51.6
14–24 (Middle) 50.8
5–13 (Early) 49.5
Figure 3: Pass rates at $2\times 32\times 600$ for InternLM2.5-StepProver on miniF2F.
Figure 4: Layer-wise ablation results.
These findings suggest that steering vectors primarily affect later-stage reasoning circuits. To ensure robustness, we also compare against random steering vectors, confirming that observed improvements arise from semantically meaningful directions rather than incidental noise.
5.3 Q3: Proof Structure and Search Efficiency
Table 3 shows that steering primarily benefits shorter proofs (<<5 steps), while gains diminish for longer ones (>>15 steps). Qualitative analysis suggests that informal reasoning introduces noisy intermediate steps for long proofs, reducing effectiveness. Interestingly, steering sometimes shortens proof length by guiding the model toward alternate proof strategies.
Proof Length Without Steering With Steering (total)
<<5 76 83 (94)
6–10 16 13 (19)
10–15 10 11 (20)
15–30 15 13 (18)
>>30 11 8 (11)
Table 3: Proof length analysis for InternLM2.5-StepProver on miniF2F.
While new proofs increase substantially, so do failure cases, likely reflecting inductive biases encoded in steering vectors. Characterizing these biases will be crucial for extending steering to broader proof distributions.
5.4 Q4: Search Budget
Figure 5 shows that steering yields consistent improvements across all search budgets $N$. Gains grow with larger budgets: at $N=100$, steering triples performance (6 vs. 2 proofs), and at $N=600$, steering delivers a 38% improvement (162 vs. 117 proofs). These results suggest that steering vectors not only boost baseline performance but also scale effectively with additional computational resources by guiding the search toward more productive proof trajectories.
Figure 5: Pass rates on miniF2F with varying search budgets.
5.5 Q5: Effectiveness vs. LoRA Fine-Tuning
We compare steering against LoRA fine-tuning (Hu et al., 2021) on InternLM2. LoRA models were trained on $p^{+}$ prompts with (rank=8, $\alpha_{\text{LoRA}}$=16) and (rank=32, $\alpha_{\text{LoRA}}$=64). Performance is reported on miniF2F after each epoch (Figure 6).
LoRA with higher rank outperforms steering, but lower-rank adaptations underperform, highlighting their limited representational capacity. In contrast, steering achieves competitive performance immediately, without additional training or parameters. This demonstrates that activation steering offers a highly parameter-efficient alternative, complementing but not replacing fine-tuning.
Figure 6: Comparison of steering and LoRA fine-tuning.
5.6 Q6: Generalization Across Provers
We evaluate if the steering vector is generalizable across different provers. In particular we use Putnambench (Tsoukalas et al., 2024) for evaluating cross lingual transfer for steering vectors. A particularly compelling finding emerges from our cross-system evaluation: despite the steering vectors being derived from Lean-based datasets, the augmented model successfully proves one Rocq problem an improvement over the base model’s zero success rate. Specifically, InternLM2.5 with steering successfully constructs a proof for Putnam 1988 B1 in Rocq, while the same model fails to complete the corresponding proof in Lean. In Appendix A.3 we showcase the proof written by InternLM2.5-StepProver.
This preliminary cross-system success suggests that steering vectors may capture reasoning patterns transferable across systems. The transferability suggests that our approach enables model to capture deeper mathematical concepts that transcend the specific formal language, encoding universal reasoning strategies that benefit theorem proving across different proof assistants. This finding has important implications for developing general-purpose mathematical reasoning systems, indicating that steering vectors trained on one formal system may provide value across the broader ecosystem of interactive theorem provers.
6 Conclusion
Our work presents a deep analysis of the underlying mechanisms that drive natural language thoughts in Large Language Models for formal theorem proving. By isolating and manipulating specific activation directions associated with natural language “thoughts”, we have demonstrated a method to enhance LLMs’ capabilities in mathematical theorem proving without requiring costly fine-tuning. Our experiments showcase how language models comprehend natural language very differently compared to formal proof steps and theorems. We explore the use of activation vectors to represent the task of producing NL information (or thought) with the proof step. And it performs consistently better than base models trained for theorem proving. We provide a closer look into how language models represent theorems in activation space to generate proofs. The use of steering vectors to isolate layers responsible for formal reasoning shows the promise in this approach. We believe this provides insight into the challenges LLMs face on Olympiad-level problems and how activation steering may mitigate them.
7 Limitations and Future Work
This work investigates the use of activation steering to interpret and enhance the performance of language models on theorem-proving tasks. While our explorations are thorough and yield consistent empirical results, there are several limitations worth noting.
First, our study focuses solely on incorporating natural language information to guide the theorem-proving process. However, we do not systematically evaluate the relevance or factual correctness of the natural language inputs with respect to the underlying proof obligations. In certain cases, the model may generate correct tactics even when the provided natural language guidance is partially incorrect or misleading. Understanding this phenomenon requires a deeper analysis of the relationship between instruction quality and proof validity, which we leave for future work.
Appendix A Appendix
In this section, we present a series of representative proof trajectories illustrating both the capabilities and failure modes of language models, and demonstrate how activation steering ameliorates key deficiencies. These examples drawn from Lean and Coq benchmarks reveal not only the quantitative improvements afforded by steering vectors, but also their qualitative impact on proof structure and robustness.
A.1 MiniF2F Problems Solved
We observe that the incorporation of steering vectors results in the generation of a substantial number of new proofs that are not derivable by the base language model alone. We show a detailed plot for all the problems in miniF2F in Figure 7.
Figure 7: Comparison between problems solved with and without steering.
This improvement in proof discovery is accompanied by a corresponding increase in failure cases. We hypothesize that this phenomenon arises due to biases introduced by the data used to construct the steering vectors. Specifically, the induced vector may encode inductive priors that align well with certain classes of theorems while misaligning with others, leading to erroneous proof trajectories. This trade-off highlights the importance of characterizing and formalizing the inductive biases embedded in steering vectors, particularly when generalizing across diverse proof distributions.
A.2 Examples of proofs
A.2.1 How does Steering help in the proof?
In the example below, we demonstrate how injecting a steering vector introduces the natural language information at each step that helps the LLM to write a proof more robustly. In the example below, we observe the use of high-level overview of the next step which helps in grounding the generation. We also compare it with the unsteered counterpart of the same problem and discuss the key differences in proof writing.
Example: Mathd Algebra 156 in miniF2F benchmark.
import Mathlib
import Aesop
set_option maxHeartbeats 0
open BigOperators Real Nat Topology Rat
theorem mathd_algebra_156 (x y : $\mathbb{R}$) (f g : $\mathbb{R}$ $\rightarrow$ $\mathbb{R}$) (h0 : $\forall$ t, f t = t ^ 4)
(h1 : $\forall$ t, g t = 5 * t ^ 2 - 6) (h2 : f x = g x) (h3 : f y = g y) (h4 : x ^ 2 < y ^ 2) :
y ^ 2 - x ^ 2 = 1 := by
-- Apply function definitions to rewrite f x and g x in terms of powers of x
have eq1 : x^4 = 5 * x^2 - 6 := by
rw [h0, h1] at h2
linarith
-- Apply function definitions to rewrite f y and g y in terms of powers of y
have eq2 : y^4 = 5 * y^2 - 6 := by
rw [h0, h1] at h3
linarith
-- Rearranged form: x^4 - 5 x^2 + 6 = 0
-- Use known factorization of the quartic as product of quadratics
have h5 : (x^2 - 2) * (x^2 - 3) = 0 := by
have h1 : x^4 - 5 * x^2 + 6 = 0 := by linarith
have h2 : x^4 - 5 * x^2 + 6 = (x^2 - 2) * (x^2 - 3) := by
ring
nlinarith
-- Apply same reasoning for y using identical polynomial structure
have h6 : (y^2 - 2) * (y^2 - 3) = 0 := by
have h1 : y^4 - 5 * y^2 + 6 = 0 := by linarith
have h2 : y^4 - 5 * y^2 + 6 = (y^2 - 2) * (y^2 - 3) := by
ring
nlinarith
-- Conclude that x^2 must be either 2 or 3 based on zero product property
have hx2 : x^2 = 2 $\vee$ x^2 = 3 := by
cases (mul_eq_zero.mp h5) with
| inl h =>
left
nlinarith
| inr h =>
right
nlinarith
-- Similarly, conclude that y^2 must be either 2 or 3
have hy2 : y^2 = 2 $\vee$ y^2 = 3 := by
cases (mul_eq_zero.mp h6) with
| inl h =>
left
nlinarith
| inr h =>
right
nlinarith
-- Perform case analysis on possible values of x^2 and y^2
cases hx2 with
| inl hx2l =>
-- Case: x^2 = 2
cases hy2 with
| inl hy2l =>
-- Case: y^2 = 2, contradicts h4 : x^2 < y^2
exfalso
nlinarith
| inr hy2r =>
-- Case: y^2 = 3, compute difference
have h : y^2 - x^2 = 1 := by
rw [hy2r, hx2l]
linarith
linarith
| inr hx2r =>
-- Case: x^2 = 3
cases hy2 with
| inl hy2l =>
-- Case: y^2 = 2, contradicts x^2 < y^2
exfalso
nlinarith
| inr hy2r =>
-- Case: y^2 = 3, contradicts x^2 < y^2
exfalso
nlinarith
Comparison with Unsteered Proof The unsteered model demonstrates a breakdown in aligning its reasoning with the algebraic form of the goal. Specifically, it bypasses the canonical application of the zero-product property via mul_eq_zero, which is essential for decomposing polynomial equalities into tractable cases. Instead, it resorts to a general-purpose disjunction strategy using eq_or_ne, which introduces branches without constructive grounding and fails to anchor the analysis in the underlying factorization. This leads to brittle reasoning that lacks semantic traceability to the original assumptions. In contrast, the steered model adheres to a proof structure that mirrors standard mathematical practice: it identifies the factorable structure of the quartic polynomials, applies mul_eq_zero to extract meaningful cases, and maintains the logical continuity required to propagate constraints across steps. The comparison reveals how steering enhances the model’s sensitivity to algebraic invariants and encourages the use of idiomatic tactics, leading to more reliable and interpretable proofs.
Diff of proofs
diff --git a/mathd_algebra_156_internlm2-5_w_steering.lean b/mathd_algebra_156_internlm2-5_wo_steering.lean
index b79f717..7f48d26 100644
--- a/mathd_algebra_156_internlm2-5_w_steering.lean
+++ b/mathd_algebra_156_internlm2-5_wo_steering.lean
@@ -3 +2,0 @@ import Aesop
@@ -5 +3,0 @@ set_option maxHeartbeats 0
@@ -11,56 +9,47 @@ theorem mathd_algebra_156 (x y : $\mathbb{R}$) (f g : $\mathbb{R}$ $\rightarrow$ $\mathbb{R}$) (h0 : $\forall$ t, f t = t
- have eq1 : x^4 = 5 * x^2 - 6 := by
- rw [h0, h1] at h2
- linarith
- have eq2 : y^4 = 5 * y^2 - 6 := by
- rw [h0, h1] at h3
- linarith
- have h5 : (x^2 - 2) * (x^2 - 3) = 0 := by
- have h1 : x^4 - 5 * x^2 + 6 = 0 := by linarith
- have h2 : x^4 - 5 * x^2 + 6 = (x^2 - 2) * (x^2 - 3) := by
- ring
- nlinarith
- have h6 : (y^2 - 2) * (y^2 - 3) = 0 := by
- have h1 : y^4 - 5 * y^2 + 6 = 0 := by linarith
- have h2 : y^4 - 5 * y^2 + 6 = (y^2 - 2) * (y^2 - 3) := by
- ring
- nlinarith
- have hx2 : x^2 = 2 $\vee$ x^2 = 3 := by
- cases (mul_eq_zero.mp h5) with
- | inl h =>
- left
- nlinarith
- | inr h =>
- right
- nlinarith
- have hy2 : y^2 = 2 $\vee$ y^2 = 3 := by
- cases (mul_eq_zero.mp h6) with
- | inl h =>
- left
- nlinarith
- | inr h =>
- right
- nlinarith
- cases hx2 with
- | inl hx2l =>
- cases hy2 with
- | inl hy2l =>
- exfalso
- nlinarith
- | inr hy2r =>
- have h : y^2 - x^2 = 1 := by
- rw [hy2r, hx2l]
- linarith
- linarith
- | inr hx2r =>
- cases hy2 with
- | inl hy2l =>
- exfalso
- nlinarith
- | inr hy2r =>
+ have hx : x ^ 4 = 5 * x ^ 2 - 6 := by rw [← h0, h1, h2]
+ have hy : y ^ 4 = 5 * y ^ 2 - 6 := by rw [← h0, h1, h3]
+ have eq_x : x ^ 4 - 5 * x ^ 2 + 6 = 0 := by linarith [hx]
+ have eq_y : y ^ 4 - 5 * y ^ 2 + 6 = 0 := by linarith [hy]
+ have factor_x : (x ^ 2 - 2) * (x ^ 2 - 3) = 0 := by
+ calc (x ^ 2 - 2) * (x ^ 2 - 3)
+ = x ^ 4 - 3 * x ^ 2 - 2 * x ^ 2 + 6 := by ring
+ _ = x ^ 4 - 5 * x ^ 2 + 6 := by ring
+ _ = 0 := eq_x
+ have factor_y : (y ^ 2 - 2) * (y ^ 2 - 3) = 0 := by
+ calc (y ^ 2 - 2) * (y ^ 2 - 3)
+ = y ^ 4 - 3 * y ^ 2 - 2 * y ^ 2 + 6 := by ring
+ _ = y ^ 4 - 5 * y ^ 2 + 6 := by ring
+ _ = 0 := eq_y
+ have x_cases : x ^ 2 = 2 $\vee$ x ^ 2 = 3 := by
+ cases’ eq_or_ne (x ^ 2 - 2) 0 with h h
+ · left; linarith [h]
+ · right
+ have : x ^ 2 - 3 = 0 := by
+ have : (x ^ 2 - 2) $\neq$ 0 := h
+ exact eq_of_mul_eq_zero_right factor_x this
+ linarith [this]
+ have y_cases : y ^ 2 = 2 $\vee$ y ^ 2 = 3 := by
+ cases’ eq_or_ne (y ^ 2 - 2) 0 with h h
+ · left; linarith [h]
+ · right
+ have : y ^ 2 - 3 = 0 := by
+ have : (y ^ 2 - 2) $\neq$ 0 := h
+ exact eq_of_mul_eq_zero_right factor_y this
+ linarith [this]
+ cases’ x_cases with hx_2 hx_3
+ · cases’ y_cases with hy_2 hy_3
+ · exfalso
+ rw [hx_2, hy_2] at h4
+ exact lt_irrefl 2 h4
+ · rw [hy_3, hx_2]
+ norm_num
+ · cases’ y_cases with hy_2 hy_3
+ · exfalso
+ rw [hx_3, hy_2] at h4
+ norm_num at h4
@@ -68 +57,2 @@
- nlinarith
+ rw [hx_3, hy_3] at h4
+ exact lt_irrefl 3 h4
A.2.2 Does Steering introduce only useful natural language information?
We analyze the proofs and the natural language incorporation added by steering vectors, and we were surprised to observe that in few cases the natural language instructions were incorrect and the model was able to come up with the correct formal proofs. This shows the capability of choosing the correct step not only depends on the correctness of the thoughts but also at the confidence of the model at the time of generation.
Example of Mathd Algebra 313 where the intermediate step is incorrectly annotated.
The failure arises from the premature application of the simplification tactic field_simp immediately after the symbolic rewrite of $i$ as $v/z$, based on the hypothesis $v=i\cdot z$. At this stage of the proof, neither $v$ nor $z$ have been instantiated with their concrete values. Consequently, the simplification is applied to the expression $i=(i\cdot z)/z$, which reduces trivially to $i=i$ under cancellation. While syntactically valid, this reasoning is circular and fails to advance toward the proof goal, which requires computing an explicit value for $i$. Furthermore, such simplification is semantically brittle because it assumes invertibility of $z$ (i.e., $z\neq 0$), which is not yet established or unfolded in the proof context. The appropriate use of field_simp should follow substitution of the concrete definitions for both $v$ and $z$, ensuring the resulting expression is a well-defined complex fraction amenable to symbolic manipulation and numeric normalization.
Lean Proof for Mathd Algebra 313
import Mathlib
import Aesop
set_option maxHeartbeats 0
open BigOperators Real Nat Topology Rat
theorem mathd_algebra_313 (v i z : $\mathbb{C}$) (h0 : v = i * z) (h1 : v = 1 + Complex.I
) (h2 : z = 2 - Complex.I) : i = 1 / 5 + 3 / 5 * Complex.I := by
-- rewrite i in terms of v and z, to isolate i from the product
have h3 : i = v / z := by
-- replace v with i x z from hypothesis h0
rw [h0]
-- simplify division by cancelling z terms directly (incorrect)
field_simp [Complex.ext_iff, h2]
-- replace i by the fraction v/z, then rewrite v and z
rw [h3, h1, h2]
-- simplify complex division into real and imaginary parts
field_simp [Complex.ext_iff]
-- normalize numbers to explicit fractions
norm_num
A.3 Transferability of Steering Vectors
To investigate the transferability of steering vectors across formal systems, we evaluated their impact when applied to a Coq-based proof assistant. The primary objective of this evaluation was to assess whether these steering vectors merely encode syntactic preferences for generating natural language interleaved with formal proof steps, or whether they also capture reasoning strategies that generalize across distinct proof environments.
In this setting, we observed that the InternLM2.5 model, when guided by the learned steering vectors, was able to successfully construct a proof for Putnam_1988_B1 in Coq. In contrast, the same model failed to complete the corresponding proof in Lean. This discrepancy suggests that the steering vectors encapsulate transferable reasoning features that are not specific to the prover, possibly due to differences in tactic libraries, proof term verbosity, or the model’s learned prior over formal syntax. The result highlights the non-trivial interaction between the encoded reasoning direction and the target proof environment, underscoring the potential of steering vectors as cross-system inductive biases.
Problem Statement for Putnam 1988 B1
A composite number is a positive integer that can be expressed as a product (ab), where (a) and (b) are (not necessarily distinct) integers from the set ({2, 3, 4, dots}).
Problem: Show that every composite number can be written in the form
xy + xz + yz + 1,
where (x, y, z) are positive integers.
Putnam 1988 B1 proof in Coq
Require Import Arith Lia Ring.
(* A number is composite if it can be written as a product of two integers greater than 1 *)
Definition composite (n: nat) : Prop :=
exists x y, 1 < x /\ 1 < y /\ n = x * y.
Theorem Putnam_1988_B1 n :
3 < n ->
composite n ->
exists x y z : nat,
0 < x /\ 0 < y /\ 0 < z
/\ n = x*y + y*z + z*x + 1.
Proof.
(* Using the definition of composite, obtain integers a and b such that n = a * b with a, b > 1 *)
intros _ [a [b [Ha [Hb Hab]]]].
(* The goal is to express n in the form x*y + y*z + z*x + 1 where x, y, z are positive *)
(* Define x and y as one less than a and b, respectively *)
set (x := a - 1). (* Then a = x + 1 *)
set (y := b - 1). (* Then b = y + 1 *)
(* Set z to 1 in order to satisfy symmetry in the expression *)
set (z := 1).
(* Establish positivity of x, y, z *)
exists x, y, z.
(* Since a > 1, x = a - 1 > 0 *)
split; [ lia | ].
(* Since b > 1, y = b - 1 > 0 *)
split; [ lia | ].
(* z is defined as 1, so clearly positive *)
split; [ lia | ].
(* Now verify the identity n = x*y + y*z + z*x + 1 *)
unfold x, y, z.
(* Replace n by a * b as given in the hypothesis *)
rewrite Hab.
(* Simplify the right-hand side of the target expression *)
simpl.
(* Resolve the final arithmetic equivalence *)
lia.
Qed.